OpenStudy (anonymous):

Solve inequalities and describe where they should be located on a number line and if they would be opened or closed circles. 2(x + 5) > 8x – 8 –2(x – 3) ≥ 5 – (x + 3)

OpenStudy (anonymous):

-2x+3>=5-x-3 -2x+x>=-3+5-3 -x>=-1 x<=1

OpenStudy (anonymous):

what about 2(x+5) > 8x -8

OpenStudy (anonymous):

2x+10>8x-8 10+8>8x-2x 18>6x